# Poison-splat: Computation Cost Attack on 3D Gaussian Splatting

### v1

@@ C.3 Ablation study on constraint strength and attack performance
Table 6: Attack performance under different levels of constraints on NeRF Synthetic dataset. Standard deviation are reported over three individual runs.
NeRF-Synthetic Dataset
Scene Attack setting Number of Gaussians Peak GPU memory [MB] Training time [minutes]
chair clean 0.494 M $\pm$ 0.001 M 3752 MB $\pm$ 57 MB 9.45 min $\pm$ 0.55 min
$\epsilon=8/255$ 0.670 M $\pm$ 0.002 M 4520 MB $\pm$ 4 MB 11.24 min $\pm$ 0.09 min
$\epsilon=16/255$ 0.957 M $\pm$ 0.003 M 8003 MB $\pm$ 138 MB 13.41 min $\pm$ 0.44 min
$\epsilon=24/255$ 1.253 M $\pm$ 0.005 M 16394 MB $\pm$ 284 MB 15.89 min $\pm$ 0.68 min
unconstrained 4.152 M $\pm$ 0.006 M 51424 MB $\pm$ 98 MB 40.71 min $\pm$ 0.67 min
drums clean 0.385 M $\pm$ 0.001 M 3456 MB $\pm$ 47 MB 8.75 min $\pm$ 0.39 min
$\epsilon=8/255$ 0.469 M $\pm$ 0.000 M 4353 MB $\pm$ 53 MB 10.19 min $\pm$ 0.42 min
$\epsilon=16/255$ 0.695 M $\pm$ 0.002 M 7190 MB $\pm$ 261 MB 11.94 min $\pm$ 0.45 min
$\epsilon=24/255$ 0.966 M $\pm$ 0.007 M 14668 MB $\pm$ 451 MB 14.86 min $\pm$ 0.33 min
unconstrained 4.109 M $\pm$ 0.039 M 62328 MB $\pm$ 878 MB 43.25 min $\pm$ 0.75 min
ficus clean 0.265 M $\pm$ 0.000 M 3194 MB $\pm$ 42 MB 6.99 min $\pm$ 0.11 min
$\epsilon=8/255$ 0.246 M $\pm$ 0.001 M 3238 MB $\pm$ 1 MB 8.28 min $\pm$ 0.56 min
$\epsilon=16/255$ 0.273 M $\pm$ 0.001 M 3783 MB $\pm$ 51 MB 8.94 min $\pm$ 0.40 min
$\epsilon=24/255$ 0.450 M $\pm$ 0.001 M 8993 MB $\pm$ 112 MB 11.32 min $\pm$ 0.31 min
unconstrained 3.718 M $\pm$ 0.001 M 50450 MB $\pm$ 452 MB 36.92 min $\pm$ 0.86 min
hotdog clean 0.185 M $\pm$ 0.000 M 3336 MB $\pm$ 11 MB 8.57 min $\pm$ 0.30 min
$\epsilon=8/255$ 0.610 M $\pm$ 0.000 M 10362 MB $\pm$ 280 MB 12.96 min $\pm$ 0.23 min
$\epsilon=16/255$ 1.147 M $\pm$ 0.003 M 29747 MB $\pm$ 57 MB 17.43 min $\pm$ 0.61 min
$\epsilon=24/255$ 1.553 M $\pm$ 0.005 M 39087 MB $\pm$ 198 MB 21.26 min $\pm$ 0.99 min
unconstrained 4.272 M $\pm$ 0.003 M 47859 MB $\pm$ 688 MB 38.85 min $\pm$ 0.90 min
lego clean 0.341 M $\pm$ 0.001 M 3532 MB $\pm$ 91 MB 8.62 min $\pm$ 0.41 min
$\epsilon=8/255$ 0.469 M $\pm$ 0.002 M 4648 MB $\pm$ 120 MB 10.38 min $\pm$ 0.47 min
$\epsilon=16/255$ 0.805 M $\pm$ 0.001 M 9960 MB $\pm$ 132 MB 13.02 min $\pm$ 0.50 min
$\epsilon=24/255$ 1.116 M $\pm$ 0.011 M 18643 MB $\pm$ 504 MB 15.33 min $\pm$ 0.63 min
unconstrained 4.159 M $\pm$ 0.002 M 78852 MB $\pm$ 2392 MB 42.46 min $\pm$ 0.88 min
materials clean 0.169 M $\pm$ 0.001 M 3172 MB $\pm$ 32 MB 7.33 min $\pm$ 0.21 min
$\epsilon=8/255$ 0.229 M $\pm$ 0.000 M 3581 MB $\pm$ 52 MB 8.84 min $\pm$ 0.29 min
$\epsilon=16/255$ 0.410 M $\pm$ 0.001 M 4886 MB $\pm$ 66 MB 9.92 min $\pm$ 0.10 min
$\epsilon=24/255$ 0.589 M $\pm$ 0.000 M 7724 MB $\pm$ 106 MB 11.51 min $\pm$ 0.43 min
unconstrained 3.380 M $\pm$ 0.000 M 43998 MB $\pm$ 675 MB 32.40 min $\pm$ 1.02 min
mic clean 0.205 M $\pm$ 0.000 M 3499 MB $\pm$ 130 MB 8.08 min $\pm$ 0.20 min
$\epsilon=8/255$ 0.251 M $\pm$ 0.001 M 4038 MB $\pm$ 132 MB 9.29 min $\pm$ 0.03 min
$\epsilon=16/255$ 0.359 M $\pm$ 0.001 M 6437 MB $\pm$ 70 MB 10.96 min $\pm$ 0.10 min
$\epsilon=24/255$ 0.514 M $\pm$ 0.005 M 13014 MB $\pm$ 141 MB 12.89 min $\pm$ 0.52 min
unconstrained 3.940 M $\pm$ 0.004 M 61835 MB $\pm$ 586 MB 39.02 min $\pm$ 0.65 min
ship clean 0.272 M $\pm$ 0.001 M 3692 MB $\pm$ 52 MB 8.87 min $\pm$ 0.44 min
$\epsilon=8/255$ 0.516 M $\pm$ 0.003 M 5574 MB $\pm$ 55 MB 11.01 min $\pm$ 0.16 min
$\epsilon=16/255$ 1.071 M $\pm$ 0.003 M 16666 MB $\pm$ 548 MB 16.12 min $\pm$ 0.42 min
$\epsilon=24/255$ 1.365 M $\pm$ 0.002 M 29828 MB $\pm$ 192 MB 18.46 min $\pm$ 0.47 min
unconstrained 4.317 M $\pm$ 0.005 M 80956 MB $\pm$ 13 MB 44.11 min $\pm$ 0.94 min
Table 7: Attack performance under different levels of constraints on MIP-NeRF360 dataset. Standard deviation are reported over three individual runs.
MIP-NeRF360 Dataset
Scene Attack setting Number of Gaussians Peak GPU memory [MB] Training time [minutes]
bicycle clean 5.793 M $\pm$ 0.007 M 17748 MB $\pm$ 554 MB 33.42 min $\pm$ 0.22 min
$\epsilon=8/255$ 6.116 M $\pm$ 0.019 M 18608 MB $\pm$ 401 MB 34.73 min $\pm$ 0.67 min
$\epsilon=16/255$ 10.129 M $\pm$ 0.053 M 27074 MB $\pm$ 1180 MB 44.44 min $\pm$ 0.43 min
$\epsilon=24/255$ 13.265 M $\pm$ 0.080 M 34870 MB $\pm$ 1882 MB 51.37 min $\pm$ 0.26 min
unconstrained 25.268 M $\pm$ 0.627 M 63236 MB $\pm$ 537 MB 81.48 min $\pm$ 1.44 min
bonsai clean 1.294 M $\pm$ 0.002 M 10216 MB $\pm$ 21 MB 21.68 min $\pm$ 1.18 min
$\epsilon=8/255$ 1.778 M $\pm$ 0.011 M 11495 MB $\pm$ 264 MB 21.39 min $\pm$ 0.48 min
$\epsilon=16/255$ 6.150 M $\pm$ 0.063 M 20564 MB $\pm$ 58 MB 32.51 min $\pm$ 0.27 min
$\epsilon=24/255$ 9.321 M $\pm$ 0.058 M 27191 MB $\pm$ 194 MB 39.84 min $\pm$ 0.15 min
unconstrained 20.127 M $\pm$ 0.122 M 54506 MB $\pm$ 3437 MB 75.18 min $\pm$ 0.39 min
counter clean 1.195 M $\pm$ 0.005 M 10750 MB $\pm$ 104 MB 26.46 min $\pm$ 0.57 min
$\epsilon=8/255$ 1.739 M $\pm$ 0.010 M 15133 MB $\pm$ 408 MB 28.06 min $\pm$ 0.61 min
$\epsilon=16/255$ 4.628 M $\pm$ 0.014 M 29796 MB $\pm$ 187 MB 37.63 min $\pm$ 0.48 min
$\epsilon=24/255$ 6.649 M $\pm$ 0.002 M 47607 MB $\pm$ 796 MB 43.68 min $\pm$ 0.14 min
unconstrained 11.167 M $\pm$ 0.092 M 80732 MB $\pm$ 255 MB 62.04 min $\pm$ 0.13 min
flowers clean 3.508 M $\pm$ 0.008 M 12520 MB $\pm$ 263 MB 25.47 min $\pm$ 0.29 min
$\epsilon=8/255$ 3.457 M $\pm$ 0.026 M 12391 MB $\pm$ 262 MB 25.53 min $\pm$ 0.39 min
$\epsilon=16/255$ 5.177 M $\pm$ 0.008 M 16338 MB $\pm$ 566 MB 30.05 min $\pm$ 0.42 min
$\epsilon=24/255$ 7.280 M $\pm$ 0.129 M 20264 MB $\pm$ 44 MB 34.25 min $\pm$ 0.59 min
unconstrained 18.075 M $\pm$ 0.136 M 45515 MB $\pm$ 2810 MB 62.62 min $\pm$ 0.62 min
garden clean 5.684 M $\pm$ 0.014 M 17561 MB $\pm$ 194 MB 33.19 min $\pm$ 0.54 min
$\epsilon=8/255$ 5.122 M $\pm$ 0.006 M 16498 MB $\pm$ 108 MB 32.17 min $\pm$ 0.71 min
$\epsilon=16/255$ 7.697 M $\pm$ 0.016 M 23115 MB $\pm$ 505 MB 40.82 min $\pm$ 0.23 min
$\epsilon=24/255$ 10.356 M $\pm$ 0.010 M 28130 MB $\pm$ 174 MB 48.88 min $\pm$ 0.15 min
unconstrained 21.527 M $\pm$ 0.027 M 52140 MB $\pm$ 1208 MB 83.81 min $\pm$ 0.07 min
kitchen clean 1.799 M $\pm$ 0.002 M 12442 MB $\pm$ 252 MB 26.61 min $\pm$ 0.28 min
$\epsilon=8/255$ 2.675 M $\pm$ 0.001 M 13760 MB $\pm$ 117 MB 31.39 min $\pm$ 0.66 min
$\epsilon=16/255$ 6.346 M $\pm$ 0.046 M 20902 MB $\pm$ 353 MB 43.24 min $\pm$ 0.59 min
$\epsilon=24/255$ 9.445 M $\pm$ 0.040 M 27646 MB $\pm$ 614 MB 53.16 min $\pm$ 0.32 min
unconstrained 12.830 M $\pm$ 0.059 M 77141 MB $\pm$ 1444 MB 73.04 min $\pm$ 0.28 min
room clean 1.513 M $\pm$ 0.004 M 12316 MB $\pm$ 13 MB 25.36 min $\pm$ 0.09 min
$\epsilon=8/255$ 3.034 M $\pm$ 0.003 M 19542 MB $\pm$ 363 MB 32.99 min $\pm$ 0.34 min
$\epsilon=16/255$ 7.129 M $\pm$ 0.020 M 46238 MB $\pm$ 802 MB 47.98 min $\pm$ 0.72 min
$\epsilon=24/255$ 9.724 M $\pm$ 0.053 M 80142 MB $\pm$ 249 MB 56.87 min $\pm$ 0.51 min
unconstrained 16.019 M $\pm$ 0.025 M 57540 MB $\pm$ 3983 MB 76.25 min $\pm$ 0.14 min
stump clean 4.671 M $\pm$ 0.070 M 14135 MB $\pm$ 218 MB 27.06 min $\pm$ 0.31 min
$\epsilon=8/255$ 5.152 M $\pm$ 0.072 M 15342 MB $\pm$ 121 MB 28.21 min $\pm$ 0.31 min
$\epsilon=16/255$ 10.003 M $\pm$ 0.292 M 25714 MB $\pm$ 1806 MB 38.48 min $\pm$ 0.59 min
$\epsilon=24/255$ 14.292 M $\pm$ 0.050 M 34246 MB $\pm$ 1108 MB 46.89 min $\pm$ 0.49 min
unconstrained 13.550 M $\pm$ 1.257 M 36181 MB $\pm$ 3769 MB 51.51 min $\pm$ 1.90 min
Appendix D Notations

### v2

@@ D.2 Full black-box attack results
Table 6: Black-box attack results on Mip-Splatting (Yu et al., 2024) as the victim system, which can further demonstrate the robust generalization ability of our attack against unknown black-box victim systems.
Constrained Black-box Poison-splat attack with $\epsilon=16/255$ against Mip-Splatting
Scene Setting clean poisoned clean poisoned clean poisoned
NS-chair 0.372 M 1.088 M (2.92$\uparrow$) 6318 33042 (5.23$\uparrow$) 7.85 13.75 (1.75$\uparrow$)
NS-drums 0.429 M 0.773 M (1.80$\uparrow$) 7474 22036 (2.95$\uparrow$) 8.15 12.60 (1.55$\uparrow$)
NS-ficus 0.234 M 0.420 M (1.79$\uparrow$) 6144 12374 (2.01$\uparrow$) 6.35 8.97 (1.41$\uparrow$)
NS-hotdog 0.200 M 1.437 M (7.18$\uparrow$) 5350 58878 (11.01$\uparrow$) 7.08 20.08 (2.84$\uparrow$)
NS-lego 0.307 M 1.127 M (3.67$\uparrow$) 6696 48682 (7.27$\uparrow$) 7.28 16.43 (2.26$\uparrow$)
NS-materials 0.216 M 0.505 M (2.34$\uparrow$) 5278 9674 (1.83$\uparrow$) 6.72 8.88 (1.32$\uparrow$)
NS-mic 0.268 M 0.557 M (2.08$\uparrow$) 5338 19758 (3.70$\uparrow$) 9.12 12.23 (1.34$\uparrow$)
NS-ship 0.330 M 1.793 M (5.43$\uparrow$) 6500 61026 (9.39$\uparrow$) 8.75 21.05 (2.41$\uparrow$)
MIP-bicycle 8.683 M DoS 80614 DoS 47.57 DoS
MIP-bonsai 1.670 M 13.016 M (7.79$\uparrow$) 30876 80826 (2.62$\uparrow$) 23.72 61.82 (2.61$\uparrow$)
MIP-counter 1.493 M 8.329 M (5.58$\uparrow$) 19478 79904 (4.10$\uparrow$) 26.25 56.75 (2.16$\uparrow$)
MIP-flowers 3.834 M 7.281 M (1.90$\uparrow$) 50922 80286 (1.58$\uparrow$) 32.85 41.80 (1.27$\uparrow$)
MIP-garden 5.828 M 8.677 M (1.49$\uparrow$) 70446 80440 (1.14$\uparrow$) 43.20 54.43 (1.26$\uparrow$)
MIP-kitchen 2.182 M 10.734 M (4.92$\uparrow$) 37024 81006 (2.19$\uparrow$) 30.63 66.92 (2.18$\uparrow$)
MIP-room 2.080 M 12.949 M (6.23$\uparrow$) 31616 81130 (2.57$\uparrow$) 27.13 77.83 (2.87$\uparrow$)
MIP-stump 5.920 M 13.925 M (2.35$\uparrow$) 65882 80480 (1.22$\uparrow$) 34.33 56.43 (1.64$\uparrow$)
Unconstrained Black-box Poison-splat attack with $\epsilon=\infty$ against Mip-Splatting
Scene Setting clean poisoned clean poisoned clean poisoned
NS-chair 0.372 M 6.106 M (16.41$\uparrow$) 6318 80732 (12.78$\uparrow$) 7.85 57.73 (7.35$\uparrow$)
NS-drums 0.429 M 6.818 M (15.89$\uparrow$) 7474 81210 (10.87$\uparrow$) 8.15 67.67 (8.30$\uparrow$)
NS-ficus 0.234 M 4.847 M (20.71$\uparrow$) 6144 80834 (13.16$\uparrow$) 6.35 45.85 (7.22$\uparrow$)
NS-hotdog 0.200 M 6.876 M (34.38$\uparrow$) 5350 80630 (15.07$\uparrow$) 7.08 59.80 (8.45$\uparrow$)
NS-lego 0.307 M 6.472 M (21.08$\uparrow$) 6696 80668 (12.05$\uparrow$) 7.28 64.93 (8.92$\uparrow$)
NS-materials 0.216 M 5.513 M (25.52$\uparrow$) 5278 81112 (15.37$\uparrow$) 6.72 45.08 (6.71$\uparrow$)
NS-mic 0.268 M 5.433 M (20.27$\uparrow$) 5338 81020 (15.18$\uparrow$) 9.12 52.83 (5.79$\uparrow$)
NS-ship 0.330 M 6.848 M (20.75$\uparrow$) 6500 81236 (12.50$\uparrow$) 8.75 66.15 (7.56$\uparrow$)
MIP-bicycle 8.683 M DoS 80614 DoS 47.57 DoS
MIP-bonsai 1.670 M DoS 30876 DoS 23.72 DoS
MIP-counter 1.493 M DoS 19478 DoS 26.25 DoS
MIP-flowers 3.834 M DoS 50922 DoS 32.85 DoS
MIP-garden 5.828 M DoS 70446 DoS 43.20 DoS
MIP-kitchen 2.182 M DoS 37024 DoS 30.63 DoS
MIP-room 2.080 M DoS 31616 DoS 27.13 DoS
MIP-stump 5.920 M DoS 65882 DoS 34.33 DoS
Appendix E Notations
@@ Appendix F Visualizations
Appendix G Ablation Study on Perturbation Range
We compare the attack effects (in terms of number of Gaussians, GPU memory and training time) and image degradation (in terms of SSIM and PSNR compared with clean input) under different perturbation ranges $\epsilon=8/255$, $\epsilon=16/255$, $\epsilon=24/255$) on NeRF-Synthetic and MIP-NeRF360 datasets, as shown in Table 10 and Table 11.
Table 10: Attack performance (number of Gaussians, GPU memory consumption and training time) and image quality (SSIM and PSNR compared with clean images) under different levels of perturbation range on NeRF Synthetic dataset.
Scene Attack setting Number of Gaussians GPU memory Training time SSIM PSNR
chair clean 0.494 M 3752 MB 9.45 min - -
$\epsilon=8/255$ 0.670 M 4520 MB 11.24 min 0.42 34.20
$\epsilon=16/255$ 0.957 M 8003 MB 13.41 min 0.21 27.94
$\epsilon=24/255$ 1.253 M 16394 MB 15.89 min 0.16 24.35
unconstrained 4.152 M 51424 MB 40.71 min 0.07 4.63
drums clean 0.385 M 3456 MB 8.75 min - -
$\epsilon=8/255$ 0.469 M 4353 MB 10.19 min 0.45 34.05
$\epsilon=16/255$ 0.695 M 7190 MB 11.94 min 0.24 27.87
$\epsilon=24/255$ 0.966 M 14668 MB 14.86 min 0.17 24.36
unconstrained 4.109 M 62328 MB 43.25 min 0.03 4.49
ficus clean 0.265 M 3194 MB 6.99 min - -
$\epsilon=8/255$ 0.246 M 3238 MB 8.28 min 0.39 34.45
$\epsilon=16/255$ 0.273 M 3783 MB 8.94 min 0.18 28.21
$\epsilon=24/255$ 0.450 M 8993 MB 11.32 min 0.13 24.53
unconstrained 3.718 M 50450 MB 36.92 min 0.08 4.41
hotdog clean 0.185 M 3336 MB 8.57 min - -
$\epsilon=8/255$ 0.610 M 10362 MB 12.96 min 0.46 33.93
$\epsilon=16/255$ 1.147 M 29747 MB 17.43 min 0.25 27.56
$\epsilon=24/255$ 1.553 M 39087 MB 21.26 min 0.17 24.11
unconstrained 4.272 M 47859 MB 38.85 min 0.07 5.75
lego clean 0.341 M 3532 MB 8.62 min - -
$\epsilon=8/255$ 0.469 M 4648 MB 10.38 min 0.48 34.08
$\epsilon=16/255$ 0.805 M 9960 MB 13.02 min 0.28 27.88
$\epsilon=24/255$ 1.116 M 18643 MB 15.33 min 0.21 24.30
unconstrained 4.159 M 78852 MB 42.46 min 0.06 4.78
materials clean 0.169 M 3172 MB 7.33 min - -
$\epsilon=8/255$ 0.229 M 3581 MB 8.84 min 0.44 34.25
$\epsilon=16/255$ 0.410 M 4886 MB 9.92 min 0.21 27.97
$\epsilon=24/255$ 0.589 M 7724 MB 11.51 min 0.14 24.35
unconstrained 3.380 M 43998 MB 32.40 min 0.08 5.40
mic clean 0.205 M 3499 MB 8.08 min - -
$\epsilon=8/255$ 0.251 M 4038 MB 9.29 min 0.36 34.46
$\epsilon=16/255$ 0.359 M 6437 MB 10.96 min 0.15 28.17
$\epsilon=24/255$ 0.514 M 13014 MB 12.89 min 0.10 24.48
unconstrained 3.940 M 61835 MB 39.02 min 0.08 4.50
ship clean 0.272 M 3692 MB 8.87 min - -
$\epsilon=8/255$ 0.516 M 5574 MB 11.01 min 0.55 33.42
$\epsilon=16/255$ 1.071 M 16666 MB 16.12 min 0.33 27.50
$\epsilon=24/255$ 1.365 M 29828 MB 18.46 min 0.24 24.03
unconstrained 4.317 M 80956 MB 44.11 min 0.04 5.31
Table 11: Attack performance (number of Gaussians, GPU memory consumption and training time) and image quality (SSIM and PSNR compared with clean images) under different levels of constraints on MIP-NeRF360 dataset.
Scene Attack setting Number of Gaussians GPU memory (MB) Training time (min) SSIM PSNR
bicycle clean 5.793 M 17748 33.42 - -
$\epsilon=8/255$ 6.116 M 18608 34.73 0.86 30.80
$\epsilon=16/255$ 10.129 M 27074 44.44 0.67 26.24
$\epsilon=24/255$ 13.265 M 34870 51.37 0.52 23.12
unconstrained 25.268 M 63236 81.48 0.02 6.60
bonsai clean 1.294 M 10216 21.68 - -
$\epsilon=8/255$ 1.778 M 11495 21.39 0.79 32.45
$\epsilon=16/255$ 6.150 M 20564 32.51 0.51 26.83
$\epsilon=24/255$ 9.321 M 27191 39.84 0.35 23.52
unconstrained 20.127 M 54506 75.18 0.01 6.27
counter clean 1.195 M 10750 26.46 - -
$\epsilon=8/255$ 1.739 M 15133 28.06 0.80 32.24
$\epsilon=16/255$ 4.628 M 29796 37.63 0.52 26.78
$\epsilon=24/255$ 6.649 M 47607 43.68 0.35 23.45
unconstrained 11.167 M 80732 62.04 0.01 6.64
flowers clean 3.508 M 12520 25.47 - -
$\epsilon=8/255$ 3.457 M 12391 25.53 0.87 29.18
$\epsilon=16/255$ 5.177 M 16338 30.05 0.71 25.60
$\epsilon=24/255$ 7.280 M 20264 34.25 0.58 22.79
unconstrained 18.075 M 45515 62.62 0.02 6.79
garden clean 5.684 M 17561 33.19 - -
$\epsilon=8/255$ 5.122 M 16498 32.17 0.84 29.85
$\epsilon=16/255$ 7.697 M 23115 40.82 0.68 25.87
$\epsilon=24/255$ 10.356 M 28130 48.88 0.54 22.82
unconstrained 21.527 M 52140 83.81 0.04 7.43
kitchen clean 1.799 M 12442 26.61 - -
$\epsilon=8/255$ 2.675 M 13760 31.39 0.83 31.58
$\epsilon=16/255$ 6.346 M 20902 43.24 0.60 26.52
$\epsilon=24/255$ 9.445 M 27646 53.16 0.45 23.18
unconstrained 12.830 M 77141 73.04 0.03 7.76
room clean 1.513 M 12316 25.36 - -
$\epsilon=8/255$ 3.034 M 19542 32.99 0.75 32.41
$\epsilon=16/255$ 7.129 M 46238 47.98 0.45 26.64
$\epsilon=24/255$ 9.724 M 80142 56.87 0.29 23.32
unconstrained 16.019 M 57540 76.25 0.01 6.42
stump clean 4.671 M 14135 27.06 - -
$\epsilon=8/255$ 5.152 M 15342 28.21 0.84 31.20
$\epsilon=16/255$ 10.003 M 25714 38.48 0.65 25.67
$\epsilon=24/255$ 14.292 M 34246 46.89 0.49 22.98
unconstrained 13.550 M 36181 51.51 0.02 7.14
Appendix H Ablation Study on Poison Ratio
In the main paper, we conducted all other experiments with 100% poison rate, based on the assumption that attacker can fully control data collection process. This assumption is reasonable for most 3D tasks, since the datasets typically are the same scene or object captured from different angles.
To gain deeper insights into the Poison-splat attack, we explore how varying levels of data poisoning ratio can affect the attack’s effectiveness. We conduct experiments using poison ratios of 20%, 40%, 60%, and 80%, and report the number of Gaussians, GPU memory consumption, and training time in Table 12. In these experiments, the poisoned views are randomly selected. To mitigate the effects of randomness, we report the mean value and standard deviation for each setting across three individual runs.
Table 12: Attack effectiveness under different poison ratio.
Scene Poison Ratio Number of Gaussians GPU memory (MB) Training time (min)
NS-hotdog-eps16 clean 0.185 M $\pm$ 0.000 M 3336 MB $\pm$ 11 MB 8.57 min $\pm$ 0.30 min
20% 0.203 M $\pm$ 0.004 M 3286 MB $\pm$ 70 MB 8.84 min $\pm$ 0.27 min
40% 0.279 M $\pm$ 0.024 M 3896 MB $\pm$ 286 MB 9.79 min $\pm$ 0.53 min
60% 0.501 M $\pm$ 0.054 M 7367 MB $\pm$ 1200 MB 11.33 min $\pm$ 0.82 min
80% 0.806 M $\pm$ 0.018 M 13621 MB $\pm$ 861 MB 14.02 min $\pm$ 0.60 min
100% 1.147 M $\pm$ 0.003 M 29747 $\pm$ 57 MB 17.43 min $\pm$ 0.61 min
NS-ship-eps16 clean 0.272 M $\pm$ 0.001 M 3692 MB $\pm$ 52 MB 8.87 min $\pm$ 0.44 min
20% 0.321 M $\pm$ 0.004 M 3764 MB $\pm$ 70 MB 9.55 min $\pm$ 0.17 min
40% 0.472 M $\pm$ 0.021 M 4796 MB $\pm$ 293 MB 10.49 min $\pm$ 0.58 min
60% 0.718 M $\pm$ 0.003 M 7034 MB $\pm$ 407 MB 12.52 min $\pm$ 0.45 min
80% 0.924 M $\pm$ 0.004 M 11850 MB $\pm$ 510 MB 14.37 min $\pm$ 0.60 min
100% 1.071 M $\pm$ 0.003 M 16666 MB $\pm$ 548 MB 16.12 min $\pm$ 0.42 min
MIP-counter-eps16 clean 1.195 M $\pm$ 0.005 M 10750 MB $\pm$ 104 MB 26.46 min $\pm$ 0.57 min
20% 1.221 M $\pm$ 0.005 M 11043 MB $\pm$ 141 MB 27.41 min $\pm$ 0.85 min
40% 1.358 M $\pm$ 0.030 M 11535 MB $\pm$ 147 MB 27.97 min $\pm$ 0.47 min
60% 2.005 M $\pm$ 0.056 M 13167 MB $\pm$ 227 MB 28.87 min $\pm$ 0.52 min
80% 3.273 M $\pm$ 0.119 M 19578 MB $\pm$ 1417 MB 32.99 min $\pm$ 0.57 min
100% 4.628 M $\pm$ 0.014 M 29796 MB $\pm$ 187 MB 37.63 min $\pm$ 0.48 min
MIP-room-eps16 clean 1.513 M $\pm$ 0.004 M 12316 MB $\pm$ 13 MB 25.36 min $\pm$ 0.09 min
20% 1.520 M $\pm$ 0.015 M 12624 MB $\pm$ 502 MB 26.45 min $\pm$ 0.75 min
40% 1.938 M $\pm$ 0.036 M 14495 MB $\pm$ 320 MB 27.99 min $\pm$ 0.72 min
60% 3.589 M $\pm$ 0.096 M 18760 MB $\pm$ 765 MB 34.09 min $\pm$ 0.79 min
80% 5.503 M $\pm$ 0.017 M 28166 MB $\pm$ 457 MB 41.47 min $\pm$ 1.36 min
100% 7.129 M $\pm$ 0.020 M 46238 MB $\pm$ 802 MB 47.98 min $\pm$ 0.72 min
From Table 12, we have following observations:
  • Attack performance, in terms of both GPU memory usage and training time extension, becomes stronger as poisoning rate increases.
  • At higher poisoning rates (between 60%-80%), the attack exhibits greater variance due to the randomness involved in selecting which views to poison. This suggests that the choice of views to poison can also impact the effectiveness of the attack.
Appendix I Ablation Study on Defense Threshold of limiting Gaussian numbers
In the main paper, we set the defense threshold to ensure a tight wrapping of necessary Gaussians for reconstruction on the clean scene. More specifically, for three scenes we show in Section 4.4 and Figure 5, we set defense threshold as in the following Table 13:
Table 13: Number of Gaussians on clean data and as defense threshold used in Section 4.4.
Scene #Gaussians on clean data #Gaussians as defense threshold
Bicycle 5.793 Million 6 Million
Flowers 3.508 Million 4 Million
Room 1.513 Million 2 Million
For a 3DGS trainer, setting a uniform defense threshold is a notably challenging task due to the significant variance in different scenes. Implementing a low threshold can safeguard the victim’s computational resources from excessive consumption, but it may also substantially compromise the quality of reconstruction for complex scenes. Choosing the defense threshold should consider this resource-quality tradeoff. Taking room of Nerf-Synthetic as an example, we tested the defense threshold acrosss 2 million to 7 million, as shown in the following Table 14 and Figure 6:
Table 14: Defense threshold impact on GPU consumption and reconstruction PSNR.
Defense threshold Max GPU memory (MB) Reconstruction PSNR
Clean input 12316 MB 29.21
2 Million 14192 MB 19.03
3 Million 16472 MB 22.27
4 Million 18784 MB 24.11
5 Million 23642 MB 24.49
6 Million 36214 MB 28.02
Poison + No Defense 46238 MB 29.08
Figure 6: Rendering of reconstructed models under different defense threshold.
As the defense threshold becomes tighter, the maximum GPU memory usage is confined within a safer range. However, this results in a significant drop in the reconstruction’s PSNR, falling to as low as 19.03, which nearly makes it unusable. (c.f. Figure 6 (A)). What’s more, while setting a threshold at 6 million Gaussians appears adequate (achieving PSNR 28.02 dB), this same threshold yields poor rendering results for bicycle scene, with only 20.33 dB PSNR (c.f. Figure 6 (F)).
Overall, determining a universal defense threshold that preserves reconstruction quality across various scenes is challenging. Designing a more effective defense against computation cost attack remains an open question, and we hope that future research will continue to explore this area.
Appendix J Image smoothing is not an ideal defense
We will show that image smoothing as a pre-processing to Gaussian Splatting training procedure is not an ideal defense. Without a reliable detection method, defenders may only resort to universally applying image smoothing to all incoming data. This pre-processing will significantly compromise reconstruction quality.
We found that although image smoothing may reduce GPU consumption to some extent, it severly undermines efforts to preserve fine image details (Yu et al., 2024; Ye et al., 2024; Yan et al., 2024b). As illustrated in Figure 7, applying common smoothing techniques such as Gaussian filtering or Bilateral filtering (Tomasi & Manduchi, 1998) leads to a substantial degradation in reconstruction quality. For instance, on the chair scene of NeRF-Synthetic dataset, reconstruction achieves 36.91 dB PSNR without pre-processing; however, with Gaussian or Bilateral filtering, the PSNR drops sharply to around 25 dB. This level of degradation is clearly undesirable. Given these challenges, we urge the community to develop more sophiscated detection and defense mechanisms against computation cost attacks, balancing resource consumption with the preservation of image quality.
Figure 7: Using image smoothing as a pre-processing step can result in unsatisfactory model reconstruction quality, particularly regarding fine details in images. This naive defense strategy significantly degrades the PSNR by approximately 10 dB, contradicting the primary goal of the service provider to maintain high-quality outputs.
Appendix K Time efficiency of Poison-Splat Attack
According to Algorithm 1, we need to solve a bi-level optimization problem, and the time complexity is related to $T$ (inner iterations) and $\tilde{T}$ (outer iterations). However in practice, our inner iterations $T$ are far fewer than outer iterations $\tilde{T}$, and the actual time cost of attacker is even less than victim’s training time, which is an advantage of our attack.
Specifically, the total running time of algorithm 1 is decided by two parts:
  1. 1.
    Train a proxy model on clean data (Line 2 of Algorithm 1);
  2. 2.
    Solve the bi-level optimization (Line 3 to Line 13 of Algorithm 1)
For example, for unconstrained attacks on MIP-NeRF360, we set $T=6000$ and $\tilde{T}=25$, and the time of attack is totally acceptable, and is even shorter than the victim training, as shown in the Table 15 below:
Table 15: Comparison of training times for the attacker and the victim.
Scene Attacker: Proxy training + Bi-level optimization Victim: Training time
Bicycle 33.42 + 16.57 min 81.48 min
Bonsai 21.68 + 14.00 min 75.18 min
Counter 26.46 + 15.23 min 62.04 min
Flowers 25.47 + 15.05 min 62.62 min
Garden 33.19 + 15.07 min 83.81 min
Kitchen 26.61 + 14.77 min 73.04 min
Room 25.36 + 16.50 min 76.25 min
Stump 27.06 + 15.63 min 51.51 min
Lastly, the primary motivation of our work is to expose the severity of this security backdoor, and attack efficiency is beyond the scope of this paper. We hope to leave the improvement of attack efficiency in future studies.